b) For the following DE, do the following: a) Find the solution in phase - amplitude form. If we think of each equation as deii please
Answer #1

Given D e is Given conditions are x 10) = 1 and 10) = 1 2 Consider taxt loro +2D +10) <=0 fenin se The auxiliary on in tim)ventication tion . = -, +342 1 = -1 + 362 30 = 2 = C : , :: El and C = ... 5464) = ēt [ cos (34) + s sincses] : : x (o)= el